OpenStudy (anonymous):

Convert y = 5 to polar form

OpenStudy (irishboy123):

it can be very mechanical: \(x = r\ cos \ \theta\) and \(y = r \ sin \ \theta\) \(y = 5 \implies r sin \ \theta = 5\)

OpenStudy (irishboy123):

which you can finish...